Michael is the sole proprietor of a small business. In June 2018, his business income is $12,000 before consideration of any § 179 deduction. He spends $245,000 on furniture and equipment in 2018. If Michael elects to take the § 179 deduction and no bonus on a conference table that cost $25,000 (included in the $245,000 total), determine the maximum cost recovery for 2018 with respect to the conference table
Answer #1

NOL is not created under section 179 and therefore it is limited to $12000

Basis = 25000-12000=13000*14.29%(7-year MACRS)=929

Total cost recovery= 12000+929 = $12929